# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v2

@@ 7 Experiments
The experiments aim to achieve two primary goals. Firstly, we aim to empirically demonstrate the practical manifestation of distinct behaviors between AID and ITD, as outlined in the theoretical findings of Section 4. Emphasis is placed on aspects specific to the nonsmooth analysis. Secondly, we intend to evaluate the empirical performance of our stochastic method NSID presented in Algorithm 1. We implement NSID by relying on PyTorch automatic differentiation for the computation of Jacobian-vector products. For AID and ITD, we use the existing PyTorch implementations [^4] https://github.com/prolearner/hypertorch.
Refer to captionRefer to caption
@@
Elastic Net  Let $(X,y)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ be a training regression dataset. The elastic net solution $w(\lambda)$ is the minimizer of the objective function $\frac{1}{n}\lVert Xw-y\rVert^{2}+\frac{\lambda_{1}}{2}\lVert w\rVert^{2}_{2}+\lambda_{2}\lVert w\rVert_{1}$, where $\lambda=(\lambda_{1},\lambda_{2})$ are the regularization hyperparameters.
Data Poisoning  We consider a data poisoning scenario similar to the one in (Xiao et al., 2015), where an attacker would like to corrupt part of the training dataset by adding noise in order to decrease the accuracy of an Elastic-net regularized logistic regression model after training. In particular, let $c$ be the number of classes and $(\tilde{X},\tilde{y})\in\mathbb{R}^{n^{\prime}\times d}\times[c]^{n^{\prime}}$ be the examples to corrupt while $(X,y)\in\mathbb{R}^{n\times d}\times[c]^{n}$ are the clean ones. Let also $\Gamma\in\mathbb{R}^{n^{\prime}\times d}$ represent the noise and define the data poisoning elastic net solution as $w(\Gamma)=\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}f(\Gamma,w)+\frac{\lambda_{1}}{2}\lVert w\rVert_{2}+\lambda_{2}\lVert w\rVert_{1}$, where $f(\Gamma,w)=\ell(Xw,y)+\ell((\tilde{X}+\Gamma)w,\tilde{y})$ and $\ell$ is the cross-entropy loss. A strategy to find $\Gamma$ would be by approximating an element of the conservative Jacobian-vector product $D_{w}(\Gamma)^{\top}y$ where $y$ is the gradient of the cross-entropy loss on an hold out set. This setting is of particular interest, since $\Gamma$ is high dimensional and hence zero-order methods like grid or random search are less appropriate. For both settings and all considered methods, we find an approximate solution $w_{t}(\lambda)$ always by iterating the contraction map which describes the iterates of the deterministic iterative soft-thresholding algorithm (see e.g., (Combettes & Wajs, 2005)). Although this may be inefficient in the stochastic setup, it yields a fairer comparison, since both the stochastic and deterministic algorithms will have the same $w_{t}(\lambda)$ as input. Additional details are in the appendix.
@@
We consider the Elastic Net scenario and construct a synthetic supervised linear regression problem with $500$ examples and $100$ features, of which $30$ are informative. As the fixed point map $\Phi$ we use one step of iterative soft-thresholding. The appropriate choice for the step-size guarantees that $\Phi$ is a contraction, in our case we set it equal to $2/(L+\mu+2\lambda_{1})$, where $L$ and $\mu$ are the largest and smallest eigenvalues values of $n^{-1}X^{\top}X$.
@@
Refer to captionRefer to caption
Figure 2: Stochastic implicit differentiation for elastic net (left) and data poisoning (right) with constant (const) and decreasing (dec) step sizes. Mean (solid line) and the geometric standard deviation (shaded region) of the approximation error over 10 runs. SID does not converge on elastic net for a specific choice of $\lambda$ and diverges in data poisoning (hence we do not report it), while NSID converges faster (at the beginning) than the Deterministic AID-FP. Note that decreasing step-sizes provide a favorable choice.
@@ Appendix B Auxiliary Lemmas
    $\operatorname{\it{e}}(\mathcal{C}\cdot\mathcal{A},\mathcal{C}\cdot\mathcal{B})\leq\lVert\mathcal{C}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A},\mathcal{B})$ and $\operatorname{\it{e}}(\mathcal{A}\cdot\mathcal{D},\mathcal{B}\cdot\mathcal{D})\leq\lVert\mathcal{D}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{A},\mathcal{B})$
@@ Appendix D Stochastic Implicit Differentiation
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t})-B_{1}\partial_{1}T(z_{t})\rVert+\lVert B_{1}\partial_{1}T(z_{t})-B_{1}C_{1}\big{)}\rVert$
@@ E.2 Stochastic Case
Let 5.1, 5.2, E.1, E.4, E.5 hold and let $\kappa=(1-q)^{-1}$. Also, suppose that $\mathbb{E}[\lVert w_{t}(\lambda)-w(\lambda)\rVert]\leq\rho_{\lambda}(t)$, for every $t\in\mathbb{N}$. Then the output ^$\hat{\nabla}f(\lambda)$ of NSID-Bilevel (Algorithm 2) where NSID uses step sizes $\eta_{i}=\Theta(i^{-1})$ satisfies
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda),D_{f}^{\operatorname{imp}}(\lambda))^{2}]=O\left(\frac{\kappa^{5}}{k}+\kappa^{4}\left(\frac{1}{J_{2}}+\rho_{\lambda}(t)\right)+\frac{\kappa^{2}}{J_{1}}\right).$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda),D_{f}^{\operatorname{imp}}(\lambda))^{2}]=O\big{(}\kappa^{2}J_{1}^{-1}+\kappa^{5}(k^{-1}+J_{2}^{-1}+\kappa^{\alpha}t^{-1})\big{)}.$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda),D_{f}^{\operatorname{imp}}(\lambda))^{2}]=O(\kappa^{5+\alpha}t^{-1})$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}$ $f(\lambda),D_{f}^{\operatorname{imp}}(\lambda))^{2}]$
@@ Appendix F Experimental Details
$G(u,\lambda)=\operatorname{Prox}_{\eta_{\lambda}g_{\lambda}}(u)\qquad T(w,\lambda)=u-\eta_{\lambda}\nabla f_{\lambda}(u).$
In particualr, since we consider Elastic net, $\operatorname{Prox}$ is the soft-thresholding. In particular, in the case of elastc net with parameters $\lambda_{1},\lambda_{2}$ we set $\eta_{\lambda}=2/(L+\mu+2\lambda_{2})$, where $L,\mu$ are the largest and smallest eigenvalues of $X^{\top}X$, where $X$ is the design matrix of the training set. Since this theoretical estimate is too conservative for data-poisoning we set $\eta_{\lambda}=20/(L+\mu+2\lambda_{2})$, $10$ times the optimal theoretical value, and we set $\mu=0$ since we are dealing with the cross-entropy loss.
@@ F.3 Details for the AID and ITD Experiments
We construct the synthetic dataset by sampling each element of the matrix $X\in\mathbb{R}^{n\times d}$ and the vector $w$ from a normal distribution. Subsequently, we set the non-informative features of $w$ to zero and we compute the vector $y$ as $y=Xw+\epsilon$, where $\epsilon_{i}$ is Gaussian noise with mean $0.1$ and unit variance. For this experiment we set $n=500$ and $p=100$ of which $30$ are informative.
@@ F.4 Details for the Stochastic Experiments
For elastic net, we enhance the setup used for the deterministic methods by sampling the population covariance matrix randomly for the informative features. To do so, we first sample a matrix $A_{1}$ from a standard normal, then we normalize all eigenvalues by diving all of them by their maximum obtaining $A_{2}$, finally we use the normalized $A_{2}^{\top}A_{2}$ as the covariance matrix of a Gaussian distribution for the informative features. This introduces correlations among the features, thereby increasing the complexity of the problem. We also increase the number of training points from $500$ to $10K$.
For the data poisoning setup we use the MNIST dataset. We split the MNIST original train set into $30K$ example for training and $30K$ examples for validation. Additionally, we perform a random split of the training set into $X\in\mathbb{R}^{n\times d}$ and $\tilde{X}\in\mathbb{R}^{n^{\prime}\times p}$, with $p=784$ representing the number of features for MNIST images. Notably, $n^{\prime}=9K$ denotes the number of corrupted examples. It is essential to highlight that $\Gamma\in\mathbb{R}^{n^{\prime}\times p}$ and $n^{\prime}p$ is approximately $7$ million, posing a significant challenge for derivative estimation using zero-order methods. For data poisoning, we observed that the theoretical value for the step size was too conservative and hence we multiply it by $10$, to have improved convergence. We set the regularization parameters $\lambda=(0.02,0.1)$ since with this setup, the final uncorrputed linear model achieves a validation accuracy of around $80\%$ with around $90\%$ of components set to zero. We note that NSID and SID require to choose the step sizes $(\eta_{i})$, which we found to be difficult, since the theoretical values are often conservative estimates for this problem. We try two policies: constant and decreasing (as $\Theta(1/i)$) step sizes, indicated with “const” and “dec” after the method name respectively. Note that only when the step sizes are decreasing NSID is guaranteed to converge. To simplify the setup, we always set them equal at $i=0$. Moreover, we set the step size of SID equal to that of NSID, when they use the same step sizes policy.
More speifically, we set $\eta_{i}=a_{1}/(a_{2}+i)$ for NSID dec and $\eta=a_{1}/a_{2}$ for NSID const, where $a_{1}=b_{1}\beta$ and $a_{2}=b_{2}\beta$, where beta is set to the theretical value suggested in Lemma D.3 ($2/(1-q^{2})$). We tuned $a_{1},a_{2}$ manually for each setting. In particular we set $a_{1}=0.5,a_{2}=2$ for the synthetic Elastic net experiment and $a_{1}=2,a_{2}=0.01$ for Data poisoning.

### v3

@@ 7 Experiments
The experiments aim to achieve two primary goals. Firstly, we aim to empirically demonstrate the practical manifestation of distinct behaviors between AID and ITD, as outlined in the theoretical findings of Section 4. Emphasis is placed on aspects specific to the nonsmooth analysis. Secondly, we intend to evaluate the empirical performance of our stochastic method NSID presented in Algorithm 1. We implement NSID by relying on PyTorch automatic differentiation for the computation of Jacobian-vector products. For AID and ITD, we use the existing PyTorch implementations [^4] https://github.com/prolearner/hypertorch. We provide the code to reproduce our experiments at https://github.com/prolearner/nonsmooth_implicit_diff
@@
Elastic Net  Let $(X,y)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ be a training regression dataset. The elastic net solution $w(\lambda)$ is the minimizer of the objective function $\frac{1}{n}\lVert Xw-y\rVert^{2}+\lambda_{1}\lVert w\rVert_{1}+\frac{\lambda_{2}}{2}\lVert w\rVert^{2}_{2}$, where $\lambda=(\lambda_{1},\lambda_{2})$ are the regularization hyperparameters.
Data Poisoning  We consider a data poisoning scenario similar to the one in (Xiao et al., 2015), where an attacker would like to corrupt part of the training dataset by adding noise in order to decrease the accuracy of an elastic-net regularized logistic regression model after training. In particular, let $c$ be the number of classes and $(\tilde{X},\tilde{y})\in\mathbb{R}^{n^{\prime}\times d}\times[c]^{n^{\prime}}$ be the examples to corrupt while $(X,y)\in\mathbb{R}^{n\times d}\times[c]^{n}$ are the clean ones. Let also $\Gamma\in\mathbb{R}^{n^{\prime}\times d}$ represent the noise and define the data poisoning elastic net solution as $w(\Gamma)=\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}f(\Gamma,w)+\lambda_{1}\lVert w\rVert_{1}+\frac{\lambda_{2}}{2}\lVert w\rVert^{2}_{2}$, where $f(\Gamma,w)=\ell(Xw,y)/2+\ell((\tilde{X}+\Gamma)w,\tilde{y})/2$ and $\ell$ is the cross-entropy loss. A strategy to find $\Gamma$ would be by approximating an element of the conservative Jacobian-vector product $D_{w}(\Gamma)^{\top}y$ where $y$ is the gradient of the cross-entropy loss on an hold out set. This setting is of particular interest, since $\Gamma$ is high dimensional and hence zero-order methods like grid or random search are less appropriate. For both settings and all considered methods, we find an approximate solution $w_{t}(\lambda)$ always by iterating the contraction map which describes the iterates of the deterministic iterative soft-thresholding algorithm (see e.g., (Combettes & Wajs, 2005)). Although this may be inefficient in the stochastic setup, it yields a fairer comparison, since both the stochastic and deterministic algorithms will have the same $w_{t}(\lambda)$ as input. Additional details are in the appendix.
@@
We consider the Elastic Net scenario and construct a synthetic supervised linear regression problem with $100$ training examples and $100$ features, of which $30$ are informative. As the fixed point map $\Phi$ we use one step of iterative soft-thresholding. The appropriate choice for the step-size guarantees that $\Phi$ is a contraction, in our case we set it equal to $2/(L+\mu+2\lambda_{2})$, where $L$ and $\mu$ are the largest and smallest eigenvalues values of $n^{-1}X^{\top}X$.
@@
Figure 2: Stochastic implicit differentiation for elastic net (left) and data poisoning (right) with constant (const) and decreasing (dec) step sizes. Mean (solid line) and the geometric standard deviation (shaded region) of the approximation error over 10 runs. SID does not converge on elastic net for this specific choice of $\lambda$ and diverges in data poisoning (hence we do not report it), while NSID converges faster (at the beginning) than the deterministic AID-FP. Note that decreasing step-sizes provide a favorable choice.
@@ 8 Conclusions
Acknowledgements
This work was supported in part from the PNRR MUR Project PE000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)“, funded by the European Union – NextGenerationEU, and EU Project ELSA under grant agreement No. 101070617.
@@ Appendix B Auxiliary Lemmas
[…]
@@ Appendix D Stochastic Implicit Differentiation
$\leq\lVert v_{k}-\bar{v}\rVert+\frac{\lVert y\rVert}{(1-q)^{2}}\Big{[}\lVert\partial_{1}G(\bar{T}(z_{t}),\lambda)\partial_{1}T(z_{t})-B_{1}\partial_{1}T(z_{t})\rVert+\lVert B_{1}\partial_{1}T(z_{t})-B_{1}C_{1}\big{)}\rVert\Big{]}$
@@ E.2 Stochastic Case
Let 5.1, 5.2, E.1, E.4, E.5 hold and let $\kappa=(1-q)^{-1}$. Also, suppose that $\mathbb{E}[\lVert w_{t}(\lambda)-w(\lambda)\rVert]\leq\rho_{\lambda}(t)$, for every $t\in\mathbb{N}$. Then the output ^$\hat{\nabla}f(\lambda)^{\top}$ of NSID-Bilevel (Algorithm 2) where NSID uses step sizes $\eta_{i}=\Theta(i^{-1})$ satisfies
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda)^{\top},D_{f}^{\operatorname{imp}}(\lambda))^{2}]=O\left(\frac{\kappa^{5}}{k}+\kappa^{4}\left(\frac{1}{J_{2}}+\rho_{\lambda}(t)\right)+\frac{\kappa^{2}}{J_{1}}\right).$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda)^{\top},D_{f}^{\operatorname{imp}}(\lambda))^{2}]=O\big{(}\kappa^{2}J_{1}^{-1}+\kappa^{5}(k^{-1}+J_{2}^{-1}+\kappa^{\alpha}t^{-1})\big{)}.$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda)^{\top},D_{f}^{\operatorname{imp}}(\lambda))^{2}]=O(\kappa^{5+\alpha}t^{-1})$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}$ $f(\lambda)^{\top},D_{f}^{\operatorname{imp}}(\lambda))^{2}]$
@@ Appendix F Experimental Details
$G(u,\lambda)=\operatorname{Prox}_{\eta_{\lambda}g_{\lambda}}(u)\qquad T(w,\lambda)=u-\eta_{\lambda}\nabla f_{\lambda}(u)$
In particular, since in our case $g_{\lambda}$ is always the an elastic net regularization, $\operatorname{Prox}$ is the soft-thresholding and we set
$\eta_{\lambda}=\frac{2}{c(L+\mu)+2\lambda_{2}},$
where $\lambda_{2}$ is the L2 regularization parameter and $L,\mu$ are the largest and smallest eigenvalues of $n^{-1}X^{\top}X$, where $X$ is the design matrix of the training set (which contains the corrupted points in the case of data poisoning). We set $c=1$ for the elastic net experiments with synthetic data, while we set $c=0.1$ for the poisoning experiments. The choice $c=1$ yields the optimal theoretical value for the step-size when using the square loss, while we used a difference choice for data poisoning (using the cross-entropy loss) since we found the optimal theoretical value for the step-size too conservative. To set the stochastic step-size schedules we also set the (estimated) contraction constant as $q=\max\{\lvert 1-\eta_{\lambda}(cL+\lambda_{2})\rvert,\lvert 1-\eta_{\lambda}(c\mu+\lambda_{2})\rvert\}$.
@@ F.3 Details for the AID and ITD Experiments
We construct the synthetic dataset by sampling each element of the matrix $X\in\mathbb{R}^{n\times d}$ and the vector $w$ from a normal distribution. Subsequently, we set the non-informative features of $w$ to zero and we compute the vector $y$ as $y=Xw+\epsilon$, where $\epsilon_{i}$ is Gaussian noise with mean $0.1$ and unit variance. For this experiment we set $n=100$ and $p=100$ of which $30$ are informative. We use $200$ hold-out examples for the validation set.
@@ F.4 Details for the Stochastic Experiments
We start by noting that for each point in the plots for the stochastic experiments in Figure 2 $w_{t}(\lambda)$ is fixed as the last iterate of deterministic iterative soft-thresholding, so that the focus lies entirely on the computation of the derivative.
For elastic net, we enhance the setup used for the deterministic methods by sampling the population covariance matrix randomly for the informative features. To do so, we first sample a matrix $A_{1}$ from a standard normal, then we normalize all eigenvalues by diving all of them by their maximum obtaining $A_{2}$, finally we use the normalized $A_{2}^{\top}A_{2}$ as the covariance matrix of a Gaussian distribution for the informative features. This introduces correlations among the features, thereby increasing the complexity of the problem. We also increase the number of training points from $100$ to $10K$ and the number of validation points from $200$ to $20K$.
For the data poisoning setup we use the MNIST dataset. We split the MNIST original train set into $30K$ example for training and $30K$ examples for validation. Additionally, we perform a random split of the training set into $X\in\mathbb{R}^{n\times d}$ and $\tilde{X}\in\mathbb{R}^{n^{\prime}\times p}$, with $p=784$ representing the number of features for MNIST images. Notably, $n^{\prime}=9K$ denotes the number of corrupted examples. It is essential to highlight that $\Gamma\in\mathbb{R}^{n^{\prime}\times p}$ and $n^{\prime}p$ is approximately $7$ million, posing a significant challenge for derivative estimation using zero-order methods. We set the regularization parameters $\lambda=(0.02,0.1)$ since with this setup, the final uncorrputed linear model achieves a validation accuracy of around $80\%$ with around $90\%$ of components set to zero.
We note that NSID and SID require to choose the step sizes $(\eta_{i})$, which we found to be difficult, since the theoretical values are often conservative estimates for this problem. We try two policies: constant and decreasing (as $\Theta(1/i)$) step sizes, indicated with “const” and “dec” after the method name respectively. Note that only when the step sizes are decreasing NSID is guaranteed to converge. To simplify the setup, we always have the same $\eta_{0}$ for both constant and decreasing step-size policies. Moreover, we set the step size of SID equal to that of NSID, when they use the same step sizes policy. More speifically, we set $\eta_{i}=a_{1}/(a_{2}+i)$ for (N)SID dec and $\eta=a_{1}/a_{2}$ for (N)SID const, where $a_{1}=b_{1}\beta$ and $a_{2}=b_{2}\beta$, where beta is set to the theoretical value suggested in Lemma D.3 ($2/(1-q^{2})$). We tuned $a_{1},a_{2}$ manually for each setting. In particular we set $a_{1}=0.5,a_{2}=2$ for the synthetic Elastic net experiment and $a_{1}=2,a_{2}=0.01$ for Data poisoning.